1. When visiting the Italian town of Pisa, young physics student Isaac Galilei drops an apple with a weight of 2.0 N from the city's famous leaning tower. The apple falls 56m to the ground. First question: Ignoring air resistance and other subtle effects, how long does it take the apple to fall to the ground? The apple takes _____ s to fall to the ground.
1
Expert's answer
2020-04-20T19:20:56-0400

As per the given question,

weight of the apple =2N

height fallen by apple = 56m

We know that gravitational acceleration of the earth "=9.8 m\/sec^2"

"h=ut+\dfrac{gt^2}{2}"

Initially , u =0

"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\times 56}{9.8}}=\sqrt{11.42}sec"

"t=3.38 sec"

gravitational acceleration is independent of the object's mass.